# RECAST: Strengthening LLMs’ Complex Instruction Following with Constraint-Verifiable Data

### v3

@@ 5.1 Setups
Evaluation Benchmarks.
We constructed RECAST-Test, a hierarchical evaluation benchmark based on RECAST-30K. This benchmark includes four different difficulty levels, categorized by increasing constraint complexity. Each level comprises instructions with progressively more constraints, enabling fine-grained performance evaluation across tasks of varying complexity. The detailed construction method and statistical characteristics of each difficulty level are provided in Appendix A.2. To assess the generalization and general capabilities of RECAST beyond our constructed benchmark, we further evaluated it on six benchmarks, the details of these benchmarks can be found in Appendix C.2.
Evaluation Metrics.
For evaluation metrics, we use the Hard Constraint Satisfaction Rate (HSR) as the primary metric, quantifying the model’s ability to satisfy all specified constraints in an instruction simultaneously. The HSR is calculated as follows:
@@
Additionally, we define the following sub-metrics for different constraint types: (1) Rule-based Constraint Satisfaction Rate (RSR), which measures the HSR score for constraints evaluated using rule-based methods; (2) Model-based Constraint Satisfaction Rate (MSR), which measures the HSR score for constraints evaluated using LLMs; and (3) Overall Constraint Satisfaction Rate (OSR), which quantifies the HSR of all constraints, both rule-based and model-based, that are successfully satisfied simultaneously.
For the FollowBench evaluation, we similarly adopt the Hard Constraint Satisfaction Rate defined in the original benchmark. For all LLM-based evaluations, unless otherwise specified, we default to using GPT-4o Achiam et al. (2023) for assessment.
Baselines.
For our baseline comparisons, we primarily select eight high-quality, open-source complex instruction fine-tuning datasets. A detailed description of the baselines can be found in Appendix C.1. Additionally, we evaluated the performance of some of the leading large models on the benchmark we constructed.
Settings.
For all experiments, we choose Qwen2.5-7B Yang et al. (2024) and Llama-3.1-8B Grattafiori et al. (2024) as the base models. The same experimental setup is used for SFT across all datasets. We use the model after SFT as the base model of RLVC for training. Specific experimental details can be found in Appendix C.
5.2 Main Results
5.2.1 Evaluation on RECAST-Test
@@
Performance Degradation with Increasing Constraint Complexity.
A consistent performance degradation is observed across all models as constraint complexity increases from Level 1 to Level 4. This systematic decline highlights the fundamental challenge LLMs face when processing and satisfying multiple constraints simultaneously. Notably, rule-based constraints present greater difficulties than model-based constraints across all model classes. Among tested models, Gemini-2.5-Pro demonstrates superior constraint handling capabilities with a 39.7539.75% average satisfaction rate.
Effectiveness of RECAST-30K for Instruction Following.
Models fine-tuned on RECAST-30K demonstrate substantial improvements in constraint satisfaction capabilities. The RECAST-30K-SFT variant achieves a 31.2531.25% average satisfaction rate on Qwen-2.5-7B, significantly outperforming both instruction-tuned models and alternative fine-tuning approaches. This enhancement underscores the effectiveness of our constraint-focused dataset in developing robust instruction-following capabilities, particularly for navigating complex, multi-constraint scenarios.
Performance Enhancement Through RLVC Optimization.
The application of RLVC further enhances model performance across all difficulty levels. RECAST-30K-RLVC variants consistently outperform their SFT counterparts, achieving a 32.3332.33% average satisfaction rate on Qwen-2.5-7B. This improvement is particularly pronounced in higher difficulty levels (Levels 3-4), where constraint interactions become more complex. These gains confirm that our constraint-specific reinforcement learning approach effectively optimizes for simultaneous satisfaction of multiple requirements, addressing a key challenge in practical instruction following.
@@
Table 2: Results of multiple benchmarks across different models. IFEVAL uses Prompt-Level Loose Accuracy (Pr. (L)), measuring the proportion of prompts where all instructions are followed using a loose criterion; FollowBench uses Hard Satisfaction Rate (HSR), evaluating complete constraint satisfaction across multi-level instructions; BBH, GPQA, MUSR, and MMLU-PRO use standard Accuracy (Acc), measuring the percentage of correct responses. The best result in each column is highlighted in bold, while the second-highest result is indicated with underline formatting.
@@
IFEVAL Followbench BBH GPQA MUSR MMLU-PRO IFEVAL Followbench BBH GPQA MUSR MMLU-PRO
Pr. (L) HSR Acc Acc Acc Acc Pr. (L) HSR Acc Acc Acc Acc
  Conifer $43.62$ $44.69$ $35.49$ $27.61$ $37.64$ $12.89$ $33.66$ $44.73$ $52.27$ $49.80$ $31.22$ $45.54$ $40.94$ $44.08$
Crab $42.14$ $41.90$ $46.85$ $32.48$ $40.26$ $32.18$ $39.30$ $44.55$ $44.50$ $49.16$ $31.89$ $48.74$ $41.72$ $43.43$
I-SHEEP $32.90$ $29.48$ $43.63$ $28.33$ $36.30$ $29.95$ $33.43$ $36.60$ $35.38$ $45.44$ $29.95$ $43.18$ $43.09$ $38.94$
MUFFIN $24.95$ $16.78$ $28.77$ $26.64$ $33.69$ $11.66$ $23.75$ $36.78$ $42.15$ $29.42$ $29.07$ $37.93$ $15.92$ $31.88$
ShareGPT $46.40$ $45.41$ $45.96$ $30.02$ $40.16$ $30.09$ $39.67$ $47.13$ $49.06$ $49.00$ $31.20$ $44.95$ $40.13$ $43.58$
Evol-Instruct $42.70$ $36.60$ $44.31$ $28.41$ $41.75$ $28.03$ $36.97$ $45.29$ $46.34$ $51.33$ $32.55$ $47.41$ $43.18$ $44.35$
Suri $18.30$ $4.02$ $29.50$ $26.09$ $33.40$ $11.81$ $20.52$ $12.75$ $5.59$ $29.23$ $26.60$ $41.79$ $11.69$ $21.28$
Tülu 3 Persona IF $73.94$ $55.44$ $30.10$ $28.64$ $38.68$ $12.50$ $39.88$ $73.94$ $56.94$ $44.75$ $31.06$ $44.13$ $41.32$ $48.69$
RECAST-30K-SFT $76.34$ $57.10$ $29.85$ $26.77$ $37.39$ $11.72$ $39.86$ $73.38$ $59.60$ $38.87$ $30.02$ $39.91$ $37.87$ $46.61$
RECAST-30K-RLVC $77.39$ $61.76$ $30.03$ $30.81$ $38.47$ $12.12$ $41.76$ $74.01$ $63.23$ $34.69$ $29.84$ $46.94$ $37.07$ $47.63$
@@ 5.2.2 General Capability Evaluation
Table 2 evaluates RECAST’s impact on general model capabilities across six diverse benchmarks. Despite training on only 30K examples focused on constraint-following, RECAST-30K-RLVC demonstrates strong performance across all methods, achieving scores of $41.76$ on Llama3.1-8B and $47.63$ on QWEN2.5-7B. The models demonstrate strong performance not only on instruction-following tasks such as IFEVAL with $77.39$ and Followbench with $61.76$, but also maintain competitive results on reasoning benchmarks, including BBH and MuSR, as well as knowledge evaluation tasks like GPQA and MMLU-PRO. This balanced improvement across fundamentally different task types confirms that RECAST’s constraint-verifiable training approach enhances models’ general capabilities without sacrificing performance on traditional benchmarks, validating its effectiveness as a comprehensive training methodology.
5.3 Ablations and Analysis
Impact of constraint type during complex instruction learning.
Impact of constraint number during complex instruction learning.
Impact of specific components of RECAST.
  Configuration Average
Llama-3.1-8B Qwen2.5-7B
  RECAST-30k $30.54$ $31.25$
Response-Only Enhancement $27.00$ $25.54$
Instruction-Only Enhancement $18.33$ $13.88$
Without RECAST $26.96$ $26.83$
To isolate component contributions, we evaluated RECAST variants with specific elements removed. Complete RECAST-30k consistently outperforms all ablated configurations. Instruction-Only Enhancement exhibits the most significant performance degradation (12.2112.21% and 17.3717.37% decreases for Llama and Qwen), revealing that constraint-augmented responses without corresponding response modifications create instruction-response inconsistencies that confuse the model, substantially hampering performance. Response-Only Enhancement performs comparably to Without RECAST, suggesting that constraint specification in instructions drives most performance gains. These results confirm that RECAST’s complete instruction-response synthesis pipeline is essential for optimal constraint satisfaction capabilities.
RLVC Training Dynamics.
Figure 5: RLVC Training Dynamics.
To analyze the effectiveness of our reinforcement learning approach, we tracked both reward metrics and constraint satisfaction performance throughout the RLVC training process. Figure 5 illustrates the reward progression. This reward optimization corresponds directly to improved constraint satisfaction capabilities. As shown in the evaluation data, the Rule-based Constraint Satisfaction Rate steadily improves from an initial 55% at step 0 to 1111% at step $200$. The reward growth (approximately 55% increase) reflects the inherent difficulty of the constraint satisfaction task. Unlike simpler RL objectives with clearer optimization landscapes, constraint satisfaction requires balancing multiple, potentially competing requirements simultaneously. This complexity creates a challenging reward space where even small improvements represent meaningful learning. These results confirm that our constraint-specific reward design effectively guides model optimization toward improved rule-based constraint handling capabilities, demonstrating the value of targeted reinforcement learning for complex instruction-following tasks.
@@ 6 Conclusion
In conclusion, we presented RECAST, a novel framework that systematically constructs training data with verifiable constraints to enhance large language models’ complex instruction-following capabilities, along with RLVC, a reinforcement learning approach that provides fine-grained feedback through constraint-specific rewards. Our comprehensive experiments demonstrated that models fine-tuned with RECAST significantly outperform corresponding instruct models across multiple benchmarks, while RLVC further enhances performance by effectively addressing limitations of traditional single-reward approaches. Through detailed analysis, we identified key factors affecting instruction-following reliability, finding that both constraint types and quantity significantly impact model performance in complex scenarios. The verifiable nature of our approach enables precise assessment of constraint satisfaction, providing valuable insights for future research on multi-objective optimization in language models. As AI systems are increasingly deployed in scenarios requiring adherence to complex, multi-faceted instructions, our work represents an important step toward more reliable and controlled language model behavior in real-world applications.
Appendix A Dataset Analysis
@@
In this section, we provide a detailed analysis of RECAST, the dataset we constructed for improving complex instruction following, with a particular focus on the characteristics of the constraints we designed and their distribution across different types.
A.1 RECAST-30K

### v4

@@ 4.1 Setups
Evaluation Benchmarks. We constructed RECAST-Test, a hierarchical benchmark with four difficulty levels defined by increasing constraint complexity, enabling fine-grained evaluation across tasks of varying difficulty. Crucially, RECAST-Test contains substantially more constraints than prior benchmarks, providing a more rigorous assessment of LLMs’ complex instruction-following capabilities. Details of its construction and statistics are in Appendix B.4. To further assess generalization, we also evaluate RECAST on four external benchmarks (Appendix D.2).
Evaluation Metrics. For evaluation metrics, we use the Hard Constraint Satisfaction Rate (HSR) as the primary metric, quantifying the model’s ability to satisfy all specified constraints in an instruction simultaneously. The HSR is calculated as follows:
@@
Additionally, we define the following sub-metrics for different constraint types: (1) Rule-based Constraint Satisfaction Rate (RSR), which measures the HSR score for constraints evaluated using rule-based methods; (2) Model-based Constraint Satisfaction Rate (MSR), which measures the HSR score for constraints evaluated using LLMs; and (3) Overall Constraint Satisfaction Rate (OSR), which quantifies the HSR of all constraints, both rule-based and model-based, that are successfully satisfied simultaneously. For the FollowBench evaluation, we similarly adopt the Hard Constraint Satisfaction Rate defined in the original benchmark. For all LLM-based evaluations, unless otherwise specified, we default to using GPT-4o (Achiam et al., 2023) for assessment.
Baselines. For our baseline comparisons, we primarily select eight high-quality, open-source complex instruction fine-tuning datasets. A detailed description of the baselines can be found in Appendix D.1. Additionally, we evaluated the performance of some of the leading large models on the benchmark we constructed.
Settings. For all experiments, we choose Qwen2.5-7B (Yang et al., 2024) and Llama-3.1-8B (Grattafiori et al., 2024) as the base models. The same experimental setup is used for SFT across all datasets. We use the model after SFT as the base model of RLVC for training. Specific experimental details can be found in Appendix D.
@@ 4.2.1 Evaluation on RECAST-Test
Performance Degradation with Increasing Constraint Complexity. Across all models, performance consistently declines from Level 1 to Level 4, underscoring the difficulty of satisfying multiple constraints simultaneously. Rule-based constraints prove more challenging than model-based ones. Gemini-2.5-Pro shows the best performance with a 39.7539.75% average satisfaction rate. Overall, complex instruction following remains difficult for current LLMs, highlighting the need for further advancement.
Effectiveness of RECAST-30K for Instruction Following. Models fine-tuned on RECAST-30K demonstrate substantial improvements in constraint satisfaction capabilities. The RECAST-30K-SFT variant achieves a 31.2531.25% average satisfaction rate on Qwen-2.5-7B, significantly outperforming both instruction-tuned models and alternative fine-tuning approaches. This enhancement underscores the effectiveness of our constraint-focused dataset in developing robust instruction-following capabilities, particularly for navigating complex, multi-constraint scenarios.
Performance Enhancement Through RLVC Optimization. The application of RLVC further enhances model performance across all difficulty levels. RECAST-30K-RLVC variants consistently outperform their SFT counterparts, achieving a 32.3332.33% average satisfaction rate on Qwen-2.5-7B. This improvement is particularly pronounced in higher difficulty levels (Levels 3-4), where constraint interactions become more complex. These gains confirm that our constraint-specific reinforcement learning approach effectively optimizes for simultaneous satisfaction of multiple requirements, addressing a key challenge in practical complex instruction following.
@@
4.2.2 Evaluation on Additional Instruction-Following Benchmarks.
Table 2 reports results on two widely used instruction-following benchmarks, IFEVAL and FollowBench. RECAST-30K-SFT already surpasses all baseline methods, and RECAST-30K-RLVC further improves performance, achieving the best results on both Llama3.1-8B and QWEN2.5-7B. These consistent gains on out-of-domain benchmarks demonstrate that RECAST not only improves complex constraint-following but also generalizes effectively to traditional instruction-following tasks, rather than merely overfitting to the constraint patterns seen during training.
Table 2: Results on instruction-following benchmarks (IFEVAL and FollowBench). IFEVAL uses Prompt-Level Loose Accuracy (Pr. (L)), while FollowBench uses Hard Satisfaction Rate (HSR). The best result in each column is highlighted in bold, while the second-highest result is indicated with underline formatting.
@@
IFEVAL Followbench IFEVAL Followbench
Pr. (L) HSR Pr. (L) HSR
  Conifer $43.62$ $44.69$ $44.16$ $44.73$ $52.27$ $48.50$
Crab $42.14$ $41.90$ $42.02$ $44.55$ $44.50$ $44.53$
I-SHEEP $32.90$ $29.48$ $31.19$ $36.60$ $35.38$ $35.99$
MUFFIN $24.95$ $16.78$ $20.87$ $36.78$ $42.15$ $39.47$
ShareGPT $46.40$ $45.41$ $45.91$ $47.13$ $49.06$ $48.10$
Evol-Instruct $42.70$ $36.60$ $39.65$ $45.29$ $46.34$ $45.82$
Suri $18.30$ $4.02$ $11.16$ $12.75$ $5.59$ $9.17$
Tülu 3 Persona IF $73.94$ $55.44$ $64.69$ $73.94$ $56.94$ $65.44$
RECAST-30K-SFT $76.34$ $57.10$ $66.72$ $73.38$ $59.60$ $66.49$
RECAST-30K-RLVC $77.39$ $61.76$ $69.58$ $74.01$ $63.23$ $68.62$
@@ 4.2.3 General Capability Evaluation
Table 3: Results on general capability benchmarks (GPQA and MUSR). All metrics use standard Accuracy (Acc), measuring the percentage of correct responses. The best result in each column is highlighted in bold, while the second-highest result is indicated with underline formatting.
  Models Llama3.1-8B Avg. QWEN2.5-7B Avg.
GPQA MUSR GPQA MUSR
Acc Acc Acc Acc
  Conifer $27.61$ $37.64$ $32.63$ $31.22$ $45.54$ $38.38$
Crab $32.48$ $40.26$ $36.37$ $31.89$ $48.74$ $40.32$
I-SHEEP $28.33$ $36.30$ $32.32$ $29.95$ $43.18$ $36.57$
MUFFIN $26.64$ $33.69$ $30.17$ $29.07$ $37.93$ $33.50$
ShareGPT $30.02$ $40.16$ $35.09$ $31.20$ $44.95$ $38.08$
Evol-Instruct $28.41$ $41.75$ $35.08$ $32.55$ $47.41$ $39.98$
Suri $26.09$ $33.40$ $29.75$ $26.60$ $41.79$ $34.20$
Tülu 3 Persona IF $28.64$ $38.68$ $33.66$ $31.06$ $44.13$ $37.60$
RECAST-30K-SFT $26.77$ $37.39$ $32.08$ $30.02$ $39.91$ $34.97$
RECAST-30K-RLVC $30.81$ $38.47$ $34.64$ $29.84$ $46.94$ $38.39$
In addition, we evaluate RECAST on two general capability benchmarks (Table 3), where RECAST-30K-RLVC achieves average scores of $34.64$ on Llama3.1-8B and $38.39$ on QWEN2.5-7B. This indicates that while RECAST substantially improves performance on complex instruction following, it also preserves competitive performance on reasoning-oriented benchmarks like MUSR and knowledge-focused evaluations such as GPQA.
5 Related Work
Complex Instruction Datasets Construction. Recent work has proposed various methods for constructing complex instruction datasets. Xu et al. (2024) proposed Evol-Instruct, which increases difficulty through iterative rewriting. Sun et al. (2024) introduced Conifer, which generates multi-level constraints with GPT-4. An et al. (2025) proposed ULTRAIF, which decomposes and recombines instructions, and Liu et al. (2025) introduced AIR, which adds constraints via a model-judge cycle. Other approaches include MUFFIN by Lou et al. (2023), which expands input features, and CRaB by Qi et al. (2024), which infers constraints via back-translation. In contrast, RECAST mines diverse constraint types directly from seed responses and integrates explicit verification for both rule-based and model-based constraints, addressing limitations in constraint complexity, diversity, and quality assurance.
Complex Instruction Following Capabilities Improvement. Recent algorithmic innovations have significantly advanced instruction following capabilities. Dong et al. (2025) proposed AutoIF, which implements a self-dialogue generation process with execution-based verification, filtering training data through executable feedback. Wang et al. (2024a) introduced RNR, which extracts roles and rules from existing instructions to generate rule-compliant responses, while Cheng et al. (2025) proposed SPaR, which employs self-play tree search refinement through multiple rounds of model self-adversarial interaction to optimize response quality. The approach in (He et al., 2024) utilizes discriminative generation, applying a discriminator model to filter generated samples for higher-quality supervision. Unlike prior methods, we leverage RECAST-30K’s quantitative and qualitative constraints, using multi-channel rewards in RLVC to treat each constraint as a separate optimization target, significantly improving accuracy and consistency in complex instruction following.
@@ 6 Conclusion
We present RECAST, a scalable, low-cost framework for constructing complex instruction-following datasets through systematic mining and verification of rule- and model-based constraints. Using this, we release RECAST-30K, enabling smaller models (e.g., Llama3.1-8B) to outperform larger instruct-tuned models. Leveraging constraint verifiability, we propose RLVC, a reinforcement learning method using fine-grained, constraint-specific rewards to enhance performance. Experiments show RECAST-trained models excel on our hierarchical benchmark, generalize to IFEVAL and FollowBench, and remain competitive on general tasks (GPQA, MUSR). RECAST advances automated data construction by offering scalable, verifiable supervision, improving constraint diversity, complexity, and quality, and providing practical resources for more controllable LLMs, and advancing their real-world applications.
Ethics statement
This work focuses on constructing complex instruction following dataset and training methods for large language models (LLMs). All data used in RECAST-30K is derived from publicly available, non-sensitive sources and does not involve private or personally identifiable information. To mitigate risks of bias and harmful outputs, we systematically design rule-based and model-based constraints with explicit verification mechanisms, ensuring high-quality and safe training signals. While our framework substantially improves models’ ability to follow complex, multi-constraint instructions, we recognize that enhanced controllability can also be misused (e.g., for generating manipulative or deceptive content). We therefore emphasize that RECAST and RLVC are intended for advancing the reliability and transparency of LLMs in research and socially beneficial applications. We encourage responsible use and further auditing to address potential fairness, bias, and misuse concerns.
Reproducibility statement
We provide detailed descriptions to facilitate reproducibility of our work. In Section 2, we introduce the design and methodology of RECAST, with further details of the data construction pipeline presented in Appendix A. Additional implementation specifications, including API usage for data generation, are described in Appendix D.3. The supervised fine-tuning (SFT) setup, including hyperparameters and engineering resources, is documented in Appendix D.4. For reinforcement learning with verifiable constraints (RLVC), we describe the reward function design in Section 3, and provide theoretical formulations, implementation resources, and training dynamics in Appendix D.5. If accepted, we will release the code, data, and trained models on GitHub to further support reproducibility.
Appendix A Detailed Description of the RECAST Data Construction Pipeline
Figure 3: Overview of the RECAST framework. The RECAST pipeline generates complex instruction-following data through four steps: (1) seed data collection across diverse domains, (2) constraint construction with both rule-based and model-based verification methods, (3) instruction enhancement by integrating selected constraints, and (4) response synthesis ensuring constraints are satisfied.
To facilitate future research in complex instruction following and to support reproducibility, we provide a comprehensive description of the RECAST data construction pipeline in this appendix. The following exposition is presented in conjunction with Figure 3, which illustrates the pipeline workflow and should help readers better understand the individual steps. We believe this detailed methodology will serve as a valuable reference for future work on dataset construction and evaluation.
The pipeline begins with the Tülu 3 Persona IF dataset (Lambert et al., 2024) as the seed dataset, which contains instruction–response pairs. Based on these responses, we construct both rule-based and model-based constraints, which can be processed in parallel.
For rule-based constraints, we define multiple types (see Table 10), each associated with a validation function (rule-based validator). These functions determine whether a response satisfies the conditions for generating a rule-based constraint. When a condition is met, a matching template from our constraint template set (Appendix E.3) is selected to generate the corresponding constraint. For example, if a response is validated as JSON, we may select the template “Format your answer as valid ”” and instantiate it with “json” to form “Format your answer as valid ’json’” (Step 2.1 in Figure 2).
For model-based constraints, we define multiple types (see Table 11). An LLM first analyzes each response to identify applicable constraint types (Applicable Constraint Type Identification, Step 2.2 in Figure 2). Another prompt is then issued to generate specific constraints for each response (Step 2.2 in Figure 2; see prompt in Figure 22). Since some generated constraints may be spurious, we apply constraint filtering to retain only those actually satisfied by the response (Step 2.2 in Figure 2).
After constructing both rule-based and model-based constraints, we form a constraint pool for each response. An LLM then compares the original instruction with the constraint pool to identify missing constraints. These are incorporated into the instruction through rewriting, producing an Enhanced Instruction. To ensure rewrite quality, multiple LLMs are queried to generate candidate rewrites, followed by majority voting (Table 7) to select the best version.
Finally, given the enhanced instructions, we use multiple LLMs to generate responses. This procedure increases response diversity while ensuring alignment with the enhanced instructions. Majority voting is again applied to select the best candidate, yielding an Enhanced Response. The resulting RECAST-30K dataset thus consists of Enhanced Instruction–Enhanced Response pairs, systematically constructed through this pipeline.
Appendix B Dataset Analysis
@@
In this section, we first introduce the characteristics of the seed dataset Tülu 3 Persona IF (Lambert et al., 2024) to justify its selection as the basis for our work, and then provide a detailed analysis of RECAST, the dataset we constructed for improving complex instruction following, with a particular focus on the characteristics of the constraints we designed and their distribution across different types.
B.1 Seed Dataset Analysis
The seed dataset for our RECAST-30K is Tülu 3 Persona IF, a subset of Tülu 3 containing 30K prompt–response pairs. We chose this dataset as our starting point because it spans a wide range of domains and includes responses rich in content—making it especially well-suited for constructing diverse and realistic constraints in our RECAST pipeline.
To further illustrate the breadth of task coverage in our seed dataset, we provide below a breakdown of the domain distribution. This diversity is a key factor that enables our framework to generate a wide variety of instruction-following tasks with meaningful, verifiable constraints. This domain-level diversity ensures that our RECAST-30K dataset supports constraint generation across a broad spectrum of real-world scenarios.
Table 4: Domain distribution of the Tülu 3 Persona IF seed dataset.
Domain Number of Instances
Storytelling 2334
Social Media Post 740
Advertising Copy 1196
Scriptwriting & Dialogue 616
Poetry 278
Essay/Blog/Review 3390
Lyrics 93
Finance 643
Legal 612
Programming & Technology 1376
Medical & Health 917
Education 4220
Daily Life 981
Internet/IT 126
Data Analysis 997
Translation 147
Summarization 992
Information Retrieval 912
Science Popularization 1556
News Reporting 610
Entertainment 294
Business & Management 1296
Politics & International Affairs 672
Scientific Research 577
Psychology 143
Role Playing 233
Other 1819
Our pipeline method is not limited to the currently used seed dataset; it is equally applicable to other corpora containing rich information. We chose this dataset primarily for ease of engineering implementation.
B.2 RECAST-30K